# SNaRe [Uncaptioned image]: Domain-aware Data Generation for Low-Resource Event Detection

### v2

@@ Abstract
Tanmay Parekh      Yuxuan Dong      Lucas Bandarkar      Artin Kim I-Hung Hsu      Kai-Wei Chang      Nanyun Peng Computer Science Department, University of California, Los Angeles        Google {tparekh, violetpeng, kwchang}@cs.ucla.edu
@@ 2 Problem Definition
We focus on the task of Event Detection (ED) Sundheim (1992); Doddington et al. (2004) for this work. ED aims to extract mentions of any events of interest from natural language text. Following ACE 2005 Doddington et al. (2004), we define an event as something that happens or describes a change of state and is labeled by a specific event type. The word/phrase that most distinctly highlights the occurrence of the event is defined as the event trigger, and the trigger-event type pair is known as the event mention. Event Detection requires extracting the event triggers from the sentence and classifying them into one of the pre-defined event types. We provide an illustration of this task below, where arrested and campaigns trigger the events of Justice: Arrest-Jail and Conflict: Demonstrate, respectively.
@@ 3 Related Works
Figure 2: Model Architecture Diagram highlighting the various components of SNaRe. First, Scout extracts and filters domain-specific triggers, then Narrator generates passages conditioned on these triggers. Finally, Refiner adds any missing annotations and sample $N$ data points per event for downstream training.
@@
Event Detection (ED) has been studied extensively Sundheim (1992); Grishman and Sundheim (1996), leading to diverse datasets in news Doddington et al. (2004); Song et al. (2015); Ellis et al. (2015), Wikipedia Li et al. (2021); Pouran Ben Veyseh et al. (2022), and general domains Wang et al. (2020); Parekh et al. (2023), as well as niche areas like biomedical Pyysalo et al. (2012); Kim et al. (2011, 2013), multimedia Li et al. (2020), cybersecurity Satyapanich et al. (2020), epidemiology Parekh et al. (2024b, c), and pharmacovigilance Sun et al. (2022). To address the growing need for event detection across expanding domains, prior works have explored transfer learning via Abstract Meaning Representation Huang et al. (2018), Semantic Role Labeling Zhang et al. (2021), and Question Answering Lyu et al. (2021). Reformulating ED as a conditional generation task has also aided low-resource training Hsu et al. (2022, 2023b); Huang et al. (2022). Recently, LLM-based reasoning Li et al. (2023a); Gao et al. (2023); Wang et al. (2023b) and transfer-learning Cai et al. (2024) has been explored, but their performance remains inferior to supervised models Huang et al. (2024). This motivates efforts in LLM-powered synthetic data generation for low-resource ED.
@@ 4 Methodology - SNaRe
In this work, we focus on domain-aware synthetic data generation using Large Language Models (LLMs) to alleviate the need for expert-annotated training data. By generating a large dataset $D_{s}=\{(X,Y)\}$, we can train domain-specific downstream ED models using minimal supervised data.
@@
To mitigate these issues, we propose SNaRe [Uncaptioned image], a domain-aware high-quality data synthesizer, comprising three components: Scout, Narrator, and Refiner. Scout studies unlabeled target domain data $D_{T}^{\prime}$ to curate domain-specific triggers, in turn, reducing domain drift. Narrator generates domain-specific sentences conditioned on Scout’s curated triggers, while Refiner adds additional annotations to ensure high-quality labels. Overall, SNaRe is a training-free LLM inference pipeline and easily deployable and scalable across domains. We provide our architectural diagram in Figure 2 and explain each component of our pipeline below.
@@ 4.1 Scout
Thus, unlike past works Ma et al. (2024) that utilize only LLMs’ internal knowledge for trigger generation, we develop Scout, which extracts high-precision domain-specific triggers using unlabeled target domain data $D_{T}^{\prime}$. Specifically, trigger extraction involves a two-stage prompt setup: (1) The first stage is tasked with identifying and filtering possible event types mentioned in the target domain sentence, and (2) The second stage aims to find the most appropriate trigger word from the unlabeled sentence for each filtered event type. We illustrate this setup in § A.
@@ 4.3 Refiner
To account for such missing annotations, we introduce Refiner - tasked to annotate missing events in the generated sentence. Here, we simply prompt the LLM to find all event mentions in the generated sentence (illustrated in Figure 9) and append them to the original sampled $Y$. To avoid adding noisy labels, we only add annotations for newly discovered events. To further improve data quality, we apply an automated rule to remove passages that do not mention the target trigger. Additionally, we standardize trigger annotations by correcting variations in trigger word forms. Such normalization and conservative filtering further aid in ensuring high data quality. Finally, we apply a greedy sampling algorithm to sample $N=50$ instances $(X^{\prime},Y)$ per event type to create our final synthetic dataset $D_{s}$.
@@ 5 Experimental Setup
We follow STAR for the implementation of the baseline models and most hyperparameter settings. For SNaRe’s passage generation, we select the top $t=10$ triggers (except $t=8$ for GENIA) for passage generation. We generate $N=50$ datapoints per event type for each generation strategy. All our experimental results are reported over an average of three runs. Additional details are provided in Appendix C.
@@ 6.1 Zero-shot Results
On average, SNaRe beats STAR by 17.3% Eve-I F1 and 16.3% Tri-C F1 points – demonstrating how domain-specific cues from unlabeled data aid in tackling domain drift. Compared to Weak Sup, SNaRe provides average gains of 3.6% Eve-I F1 and 3.3% Tri-C F1, suggesting how cleaner label quality can help improve model performance.
@@ 6.2 Few-shot Results
We also study the various methods in the presence of small annotated data as part of our few-shot experiments. Specifically, we study the $k=2$ and $k=5$ few-shot settings, where $k$ annotated examples per event type are utilized. We utilize the $k$-shots as in-context examples in the LLM prompts and append these few-shot examples to the synthesized training data as well. Additionally, we consider another baseline (Supervised) of downstream models trained only on the $k$-shot examples. We present the Tri-C results for all the datasets for the Llama3-8B model in Figure 4. Similar to zero-shot results, we observe that SNaRe consistently beats all other baseline models. On average, SNaRe outperforms STAR and Weak Sup by 5.4% Tri-C F1 and 7% Tri-C F1 respectively.
@@ 7 Analysis
Table 3: Ablation study for SNaRe’s Scout and Refiner measured as Tri-C F1 performance across datasets.
@@ 7.1 Ablation study
Table 3 shows the ablation study for Scout and Refiner. For ablating Scout, we replace it by prompting LLM to directly generate triggers. We observe how both Scout and Refiner are critical, with average performance reductions of 3.8% F1 and 6% F1 upon removing the respective components.
@@ 7.3 Comparison with Data mixing
Data-mixing Hoffmann et al. (2022); Xie et al. (2023) is a widely used technique to leverage complementary information across datasets to promote robust downstream model training. We mix data from our two baselines of Weak Sup and STAR as a hybrid baseline to compare with SNaRe. To keep the comparisons fair, we consider $N/2=25$ data instances per event type from each dataset. Results from Table 5 demonstrate how SNaRe beats the data-mixing based hybrid model by 5-6% F1, underlining the significance of our three-stage model design over simpler data-mixing.
@@ 7.4 Analyzing domain drift and label quality
To further study the quality and relevance of the label, we performed a human evaluation. Specifically, a human expert in ED is tasked with scoring generations (between 1-5) on the naturalness of the sentence specific to the target domain, the relevance of the event to the semantic actions described in the generated sentence, and the annotation quality evaluating the selection of triggers (details in § D.3). We provide the averaged scores across the three datasets for 90 samples in Table 8.222Since Weak Sup annotates the unlabeled target domain sentences, event relevance is analogous to annotation quality and we do not explicitly evaluate it. Weak Sup has high sentence quality but poor label annotations; STAR suffers from poor event relevance indicating domain drift. Overall, SNaRe performs the best with high annotation quality and event relevance.
@@ 8 Conclusion and Future Work
Appendix A Additional details about SNaRe
We provide illustrations of the various prompts used in SNaRe. For Scout, we consider a two-prompt approach. The first prompt aims to identify whether events of interest are mentioned in the passage, as illustrated in Figure 6. This prompt comprises the task statement, full event ontology with definitions, the task instructions, and the query sentence. The second prompt identifies the trigger corresponding to the event of interest, as illustrated in Figure 7. Here, we specify the task definition, the event ontology details, and the query text with the task instructions.
For Narrator, we prompt the LLM with the task definition, the event ontology with event definitions, and the query comprising the sampled target structure. The LLM is asked to generate a sentence using the specified trigger words while mentioning the events of interest. We illustrate this prompt in Figure 8.
Finally, we provide the Refiner prompt in Figure 9. This prompt aims to identify any missing event mentions from the passage.
Appendix B Data Statistics
@@
Appendix C Implementation Details
@@ C.1 LLM-based Generation
C.2 Few-shot Implementation Details
@@
C.3 Downstream Model Training
@@ C.4 LLM Fine-tuning
Appendix D Additional analyses
@@
Figure 5: Model performance for SNaRe as keep change the number of generated datapoints $N$ using Llama3-8B-Instruct for the three datasets.
D.1 STAR with domain-specific prompt
@@
Figure 6: Prompt for stage 1 of Scout.
Figure 7: Prompt for stage 2 of Scout.
Figure 8: Prompt for Narrator.
Figure 9: Prompt for Refiner.
D.2 Impact of different number of training samples
@@
D.3 Human Evaluation Details
@@
D.4 Analysis in low-resource unlabeled data settings
@@
D.5 Additional Qualitative Examples

### v3

@@ Abstract
Tanmay Parekh Yuxuan Dong Lucas Bandarkar Artin Kim I-Hung Hsu Kai-Wei Chang Nanyun Peng Computer Science Department, University of California, Los Angeles Google {tparekh, kwchang, violetpeng}@cs.ucla.edu
@@ 2 Problem Definition
We focus on the task of Event Detection Sundheim (1992); Doddington et al. (2004) for this work. ED aims to extract mentions of any events of interest from natural language text. Following ACE 2005 Doddington et al. (2004), we define an event as something that happens or describes a change of state and is labeled by a specific event type. The word/phrase that most distinctly highlights the occurrence of the event is defined as the event trigger, and the trigger-event type pair is known as the event mention. Event Detection requires extracting the event triggers from the sentence and classifying them into one of the pre-defined event types. We provide an illustration of this task below, where arrested and campaigns trigger the events of Justice: Arrest-Jail and Conflict: Demonstrate, respectively.
@@ 3 Related Works
Figure 2: Model Architecture Diagram highlighting the various components of SNaRe. First, Scout extracts and filters domain-specific triggers, then Narrator generates passages conditioned on these triggers. Finally, Refiner adds any missing annotations and sample NN data points per event for downstream training.
@@
Event Detection (ED) has been studied extensively Sundheim (1992); Grishman and Sundheim (1996), leading to diverse datasets in news Doddington et al. (2004); Song et al. (2015); Ellis et al. (2015), Wikipedia Li et al. (2021); Pouran Ben Veyseh et al. (2022), and general domains Wang et al. (2020); Parekh et al. (2023), as well as niche areas like biomedical Pyysalo et al. (2012); Kim et al. (2011, 2013), multimedia Li et al. (2020), cybersecurity Satyapanich et al. (2020), epidemiology Parekh et al. (2024b, c), and pharmacovigilance Sun et al. (2022). To address the growing need for event detection across expanding domains, prior works have explored transfer learning via Abstract Meaning Representation Huang et al. (2018), Semantic Role Labeling Zhang et al. (2021), and Question Answering Lyu et al. (2021). Reformulating ED as a conditional generation task has also aided low-resource training Hsu et al. (2022, 2023b); Huang et al. (2022). Recently, LLM-based reasoning Li et al. (2023a); Gao et al. (2023); Wang et al. (2023b); Parekh et al. (2025a) and transfer-learning Cai et al. (2024) has been explored, but their performance remains inferior to supervised models Huang et al. (2024). This motivates efforts in LLM-powered synthetic data generation for low-resource ED. Although our work only demonstrates results on ED, we believe our work can be extended to other tasks outside information extraction as well, like question-answering Rajpurkar et al. (2016); Parekh et al. (2025b) and long-form generation Suvarna et al. (2024); Chen et al. (2020); Parekh et al. (2020).
@@ 4 Methodology - SNaRe
In this work, we focus on domain-aware synthetic data generation using LLMs to alleviate the need for expert-annotated training data. By generating a large dataset $D_{s}=\{(X,Y)\}$, we can train domain-specific downstream ED models using minimal supervised data.
@@
To mitigate these issues, we propose SNaRe [Uncaptioned image], a domain-aware data synthesizer, that generate and verifies LLM generations to ensure high-quality data Hsu et al. (2024), comprising three components: Scout, Narrator, and Refiner: Scout studies unlabeled target domain data $D_{T}^{\prime}$ to curate domain-specific triggers, in turn, reducing domain drift. Narrator generates domain-specific sentences conditioned on Scout’s curated triggers, while Refiner adds additional annotations to ensure high-quality labels. Overall, SNaRe is a training-free LLM inference pipeline and easily deployable and scalable across domains. We provide our architectural diagram in Figure 2 and explain each component of our pipeline below.
Figure 3: Prompt for stage 1 of Scout.
Figure 4: Prompt for stage 2 of Scout.
Figure 5: Prompt for Narrator.
Figure 6: Prompt for Refiner.
@@ 4.1 Scout
Thus, unlike past works Ma et al. (2024) that utilize only LLMs’ internal knowledge for trigger generation, we develop Scout, which extracts high-precision domain-specific triggers using unlabeled target domain data $D_{T}^{\prime}$. Specifically, trigger extraction involves a two-stage prompt setup: (1) The first stage is tasked with identifying and filtering possible event types mentioned in the target domain sentence, and (2) The second stage aims to find the most appropriate trigger word from the unlabeled sentence for each filtered event type. We provide the prompts for Scout in Figure 3 and  4.
@@ 4.3 Refiner
To account for such missing annotations, we introduce Refiner - tasked to annotate missing events in the generated sentence. Here, we simply prompt the LLM to find all event mentions in the generated sentence (illustrated in Figure 6) and append them to the original sampled $Y$. Since these new refined labels can be noisy, we avoid updating them for the already present events from the Scout, and only add them for newly discovered events. To further improve data quality, we apply an automated rule to remove passages that do not mention the target trigger. Additionally, we standardize trigger annotations by correcting variations in trigger word forms. Such normalization and conservative filtering further aid in ensuring high data quality. Finally, we apply a greedy sampling algorithm to sample $N=50$ instances $(X^{\prime},Y)$ per event type to create our final synthetic dataset $D_{s}$.
@@ 5 Experimental Setup
We follow STAR for the implementation of the baseline models and most hyperparameter settings. For SNaRe’s passage generation, we select the top $t=10$ triggers (except $t=8$ for GENIA) for passage generation. We generate $N=50$ datapoints per event type for each generation strategy. All our experimental results are reported over an average of three runs. Additional details are provided in Appendix B.
@@ 6.1 Zero-shot Results
On average, SNaRe outperforms STAR by 17.3% Eve-I F1 and 16.3% Tri-C F1 points – demonstrating how domain-specific cues from unlabeled data aid in tackling domain drift. Compared to Weak Sup, SNaRe provides average gains of 3.6% Eve-I F1 and 3.3% Tri-C F1, suggesting how cleaner label quality can help improve model performance.
@@ 6.2 Few-shot Results
We also study the various methods in the presence of small annotated data as part of our few-shot experiments. Specifically, we study the $k=2$ and $k=5$ few-shot settings, where $k$ annotated examples per event type are utilized. We utilize the $k$-shots as in-context examples in the LLM prompts and append these few-shot examples to the synthesized training data as well. Additionally, we consider another baseline (Supervised) of downstream models trained only on the $k$-shot examples. We present the Tri-C results for all the datasets for the Llama3-8B model in Figure 8. Similar to zero-shot results, we observe that SNaRe consistently outperforms all other baseline models. On average, SNaRe outperforms STAR and Weak Sup by 5.4% Tri-C F1 and 7% Tri-C F1 respectively.
@@ 7 Analysis
Table 3: Ablation study for SNaRe’s Scout, Narrator, and Refiner measured as Tri-C F1 performance across the three datasets.
@@ 7.1 Ablation study
Table 3 shows the ablation study for Scout and Refiner. For ablating Scout, we replace it by prompting LLM to directly generate triggers. Ablating Narrator is similar to the Weak Supbaseline, while adding additional refiner and dataset statistics to remove noisy datapoints. We observe how all the components are critical, with average performance reductions of 3.8% F1, 6.6% F1, and 6% F1 upon removing the respective components of Scout, Narrator, and Refiner.
@@ 7.3 Comparison with Data mixing
Data-mixing Hoffmann et al. (2022); Xie et al. (2023) is a widely used technique to leverage complementary information across datasets to promote robust downstream model training. We mix data from our two baselines of Weak Sup and STAR as a hybrid baseline to compare with SNaRe. To keep the comparisons fair, we consider $N/2=25$ data instances per event type from each dataset. Results from Table 5 demonstrate how SNaRe outperforms the data-mixing based hybrid model by 5-6% F1, underlining the significance of our three-stage model design over simpler data-mixing.
@@ 7.4 Analyzing domain drift and label quality
To further study the quality and relevance of the label, we performed a human evaluation. Specifically, a human expert in ED is tasked with scoring generations (between 1-5) on the naturalness of the sentence specific to the target domain, the relevance of the event to the semantic actions described in the generated sentence, and the annotation quality evaluating the selection of triggers (details in § C.6). We provide the averaged scores across the three datasets for 90 samples in Table 8.222Since Weak Sup annotates the unlabeled target domain sentences, event relevance is analogous to annotation quality and we do not explicitly evaluate it. Weak Sup has high sentence quality but poor label annotations; STAR suffers from poor event relevance indicating domain drift. Overall, SNaRe performs the best with high annotation quality and event relevance.
@@ 8 Conclusion and Future Work
Acknowledgments
We express our gratitude to Anh Mac, Po-Nien Kung, and Christina Chance for their valuable time, reviews of our work, and constructive feedback. We thank the anonymous reviewers, area chairs, and the program committee for their reviews and feedback. This work was partially supported by the National Science Foundation CAREER award #2339766 and the Amazon AGI Research Award awarded to Nanyun Peng.
@@ Acknowledgments
Appendix A Data Statistics
@@
Appendix B Implementation Details
@@ B.1 LLM-based Generation
B.2 Few-shot Implementation Details
@@
B.3 Downstream Model Training
@@ B.4 LLM Fine-tuning
Appendix C Additional analyses
@@
Method Original ACE Synthetic ACE
EI TC EI TC
LLM Training 71.5 67.4 54.3 46.8
Small Model Training 72.2 68.9 57.4 50.2
Number of ACE SPEED GENIA
Triggers EI TC EI TC EI TC
$t=5$ 55.9 48.9 43.2 29.9 32.2 26.5
$t=8$ 57.9 49.7 45.1 30.8 35.2 28.9
$t=10$ 57.4 50.2 44.6 31.5 34.8 28.1
$t=12$ 56.1 49.3 44.3 31.2 33.9 27.4
Table 17: Ablating the impact of the number of triggers for the Scout component on downstream model performance across the three datasets. EI: Event Identification F1, TC: Trigger Classification F1.
@@
Figure 9: Model performance for SNaRe as keep change the number of generated datapoints NN using Llama3-8B-Instruct for the three datasets.
Trigger Selection EI TC
Frequency Ranking (current) 57.4 50.2
Sampling 54.7 47.3
Weighted Sampling 55.2 48.1
Reranking 49.6 37.9
Minimum Count Filtering 52.4 44.8
Clustering 48.0 40.7
Table 19: Ablating downstream model performance on the trigger selection strategy for the Scout component on the ACE dataset. EI: Event Identification F1, TC: Trigger Classification F1.
C.1 STAR with domain-specific prompt
@@
C.2 Impact of different number of training samples
@@
C.3 Training LLMs with synthetic data
Since we generate synthetic data using LLMs, it would be natural to fine-tune LLMs to be better at Event Detection (ED). We conduct a small experiment to test and compare the model training for LLMs and small models. Specifically, we train a Llama3-8B-Instruct model using Low Rank Adaptation (LoRA) Hu et al. (2021) and compare it with the best performing small models like DEGREE Hsu et al. (2022) and TagPrime Hsu et al. (2023a). We conduct sets of training: (1) Original ACE training data, and (2) Synthetic data generated by SNaRe for the ACE domain. We summarize our results and findings in Table LABEL:tab:llm-training. Overall, we note how our synthetic data can be used to train LLMs and improve their performance iteratively as well. However, we would like to note how small language model still learn better and provide better scores than LLMs on this task. We would also like to note that our LLM training can be suboptimal, as we tried using the original hyperparameters and didn’t tune the hyperparameters much. We utilized Llama-Factory Zheng et al. (2024) for our LLM fine-tuning.
C.4 Ablating the design choice for number of triggers
We utilize $t=10$ for ACE and SPEED and $t=8$ for GENIA. Here, we provide additional experiments to justify this design choice. Specifically, we provide the performance of training smaller models using different number of triggers ($t$) for the Scout using Llama3-8B-Instruct LLM in Table 17. Overall, we observe that changing the number of triggers affects the performance slightly, and can be optimally chosen to improve performance. However, we note that even with different $t$, SNaRe still outperforms the other baselines.
C.5 Ablating the design choice for trigger selection
We utilize frequency-based trigger selection in the Scout, but this can lead to missing out on low-frequency, long-tail, rare triggers. Our major motive is to avoid the noisy triggers which would also have low frequency. In order to justify this design choice, we conduct experiments with various other trigger selection designs, specifically: (1) Sampling: Instead of filtering and selecting, we sample uniformly from the trigger list, (2) Weighted Sampling: We sample using the extraction frequency of the triggers as the weights, (3) Reranking: Since some triggers might not be extracted correctly, we rerank the triggers based on the number of occurrences in holdout set (basically avoiding the bias of extracting them using the LLM), (4) Minimum Count Filtering: Instead of filtering based on ranks, we simply remove all triggers with a set minimum count (as this long tail might be noisy) and sample from the remaining triggers, (5) Clustering: We use k-means clustering to form clusters of similar triggers and sample one trigger from each - thus, providing enhanced diversity. We compare these various methods for trigger selection with the Scout using Llama3-8B-Instruct LLM on the ACE dataset in Table 19. Overall, this study shows how the different methods, which introduce more noise and provide different ranges of the precision-recall balance, are worse in comparison to our existing frequency ranking method.
@@ C.5 Ablating the design choice for trigger selection
C.6 Human Evaluation Details
@@
C.7 Analysis in low-resource unlabeled data settings
@@
Method Arabic Chinese
EI TC EI TC
Inference 33.0 18.7 25.0 21.0
STAR 28.1 22.0 32.3 25.9
Weak Sup 29.4 20.2 30.2 28.4
SNaRe (ours) 38.4 31.9 39.6 29.0
Table 22: Comparing SNaRe with zero-shot inference for Qwen3-8B for the ACE dataset.
C.8 Additional zero-shot multilingual experiments
In § 6.3, we demonstrated the utility of our framework for zero-shot multilingual setting using the Llama3-8B and Llama3-70B models. Here, we additionally show results using the Qwen3-8B LLM in Table 22. We note that SNaRe outperforms the other baselines for both languages and metrics, even with the Qwen3-8B LLM.
@@ C.8 Additional zero-shot multilingual experiments
C.9 Additional Qualitative Examples